An insect start moving from origin at a constant velocity of 2i+3j for one second and then with constant velocity 2i + 2j for half second the distance of the insects from origin after these displacement is
 
a) 4m
b) 5m
c) 7m
d) 10m

Grade:11

1 Answers

Khimraj
5 years ago
coordinate after 1 sec = (2*1, 3*1) = (2,3)
coordinate after one and half sec = (2+ 2*1/2, 3 + 2*1/2) = (3,4)
So distance from origin = \sqrt{3^{2} + 4^{2}} = 5m
Hope it clears.